[figure: 9.7 cm]

THE Sn HYPERFINE FIELDS IN RMn_{6}Sn_{6} (R = Mg, Zr, Lu).

EXPERIMENTAL AND THEORETICAL STUDY

aT. Mazet, ^{b}J. Toboła and ^{a}B. Malaman aLaboratoire de la Chimie du Solide Mineral, Université Henri Poincaré, 54506 Vandoeuvre les Nancy Cedex, France

bFaculty of Physics and Nuclear Techniques, Academy of Mining and Metallurgy,

Al. Mickiewicza 30, 30-059 Kraków, Poland

Ternary RMn_{6}Sn_{6} compounds are magnetic and form hexagonal structure (P 6/mmm) involving three inequivalent Sn sublattices. The coordination and arrangement of the Mn magnetic moments in relation to the Sn sites determine local spin densities on these atoms. The ^{119}Sn Mössbauer effect measurements on the ferromagnetic MgMn_{6}Sn_{6} result in the large values of H_{hf}: 220, 300 and 340 (in kGs) on Sn1, Sn2 and Sn3 sites, respectively. Conversely, ZrMn_{6}Sn_{6} and LuMn_{6}Sn_{6} show a purely antiferromagnetic structure (at least at RT), with (- + + -) ordering of the Mn magnetic moments layers. Thus, non-zero hyperfine fields on the two Sn sites only are measured. Such behaviours are also studied by the KKR band structure calculations. The theoretical values of H_{hf} well correspond to the experimental data in the ferromagnetic MgMn_{6}Sn_{6} (240, 315, 330 kGs). In order to consider a real magnetic cell in ZrMn_{6}Sn_{6} and LuMn_{6}Sn_{6} a doubled unit cell is taken into account in the KKR computations. However, a comparison with experiment seems less straightforward, due to a helimagnetic structure appearing in these samples at low temperature. On the other hand, the calculated Mn magnetic moments by the KKR method, both in the ferromagnetic and antiferromagnetic RMn_{6}Sn_{6}, remain in excellent agreement with the neutron data.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : Thomas Mazet

Address for correspondence : Laboratoire de la Chimie du Solide Mineral, Université Henri Poincaré 54506 Vandoeuvre les Nancy Cedex, France

Email address : Thomas.Mazet@lcsm.u-nancy.fr